# Monte Carlo Physics-informed neural networks for multiscale heat conduction via phonon Boltzmann transport equation

### v2

@@ 4 Summary
In this study, we develop a Monte Carlo physics-informed neural networks (MC-PINNs) for solving the phonon Boltzmann transport equation (BTE) to model the multiscale heat conduction in solid materials. Specifically, we utilize a deep neural network to approximate the energy density function in BTE, and encode the BTE as well as the boundary/initial conditions via the automatic differentiation, in MC-PINNs. In addition, we propose a novel two-step sampling approach to address the issues like “curse of dimensionality” and the inaccurate estimate of the equilibrium distribution functions in the existing sampling strategies that are widely used in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial domains (Step I), and then we draw another number of random points in the solid angular domain (Step II). The residual points in MC-PINNs are constructed using the tensor product given the points generated at the above two steps. We conduct a series of numerical experiments to validate the accuracy of the MC-PINNs, including the steady quasi-1D/quasi-2D/3D as well as a unsteady quasi-1D heat conduction problems spanning the diffusive and ballistic regimes. The results show the the MC-PINN is capable to achieving good accuracy for the multiscale head conduction from 1D to 3D. Furthermore, the present method requires quite a small number of discrete points in the solid angular domain than in the deterministic methods for solving the BTE, especially for scenarios in the ballistic regimes. For instance, we only employ 64 points in the solid angular domain at each training step in MC-PINNs for Kn ranging from 0.1 to 10 in the 3D case. In contrast, we have $80\times 80$ discrete points in the solid angular domain of the numerical method to obtain converged results. As a result, the memory cost in the MC-PINN is about 1/6 of that in the numerical method. Further, we are able to achieve a similar computational efficiency for MC-PINNs to the numerical method in the 3D test case as we utilize the transfer learning technique in the MC-PINNs.
@@
We would like to discuss that the MC-PINN developed in this work is mesh-free in the temporal-spatial-angular domain, which is promising in real-world applications with complicated domains, for instance, the design of microdevices. In addition, the MC-PINN is capable of providing good accuracy for heat conduction ranging from diffusive to ballistic with only a small number of random discrete points in the solid angular domain at each training step. It can save lots of effort in numerical methods since the optimal discretization of the solid angular space at different regimes is generally different and should be carefully designed to ensure the convergence in these methods. Also, the memory efficacy and the flexibility of treatment for the solid angular space in the present method makes it a promising tool for modeling multiscale heat conduction in large scale problems. Finally, we note that the MC-PINNs with the two-step sampling approach can be seamlessly extended to solve kinetic equations used in other disciplines, e.g., Boltzmann-BGK model for multiscale flows, the radiation transfer equation for multiscale heat transfer, etc. We will leave these interesting topics as future studies.
@@ Appendix B Ablation study
To study the effect of the number of the points in the solid angular space at each training step on the predicted accuracy of MC-PINNs, we further conduct the following experiments based on two specific cases in Sec. 3.1, i.e., Kn $=1$ and $10$. In each case, we test 6 different numbers of the points in the solid angular space, i.e. $B_{\bm{s}}=2,4,8,16,32$ and 64 points at each training step. Also, we independently run MC-PINNs 5 times with different initializations for the hyperparameters to test the robustness of the MC-PINNs. The number of $B_{\bm{x}}$ is set to $40$, which is consistent with the case in Sec. 3.1. The predicted temperature from MC-PINNs are depicted in Fig. 15. It is observed that: (1) the computational errors in general decrease with the increase of $B_{\bm{s}}$, and (2) the relative $L_{2}$ errors between the results from MC-PINNs and the reference solutions in each case are below $1\%$ as $B_{\bm{s}}>2$. In other words, we are able to achieve good accuracy as we randomly sample more than 2 points in the solid angular domain at each training step in MC-PINNs for these two specific cases.
Figure 15: 1D film heat conduction: Relative $L_{2}$ errors of the temperature with different numbers of points in the solid angular space at each training step in PINNs. (a) $\rm{Kn}=1$, and (b) $\rm{Kn=10}$. Error bars in each figure represent the standard deviation calculated from 5 independent runs with different random seeds. The blue dashed line denotes the relative $L_{2}$ error for $1\%$.
We further present the loss histories for the representative case $\rm{Kn}=1.0$ with $B_{\bm{s}}=2$ and 16 in Fig. 16. As shown, the loss for MC-PINNs with 16 random points in the solid angular space after 50,000 training steps is about two orders smaller than that with 2 random points in the solid angular domain. Also, the fluctuations in the loss are smaller for the case with a larger number of points in the solid angular space at ach training step.

### v3

@@ 4 Summary
In this study, we develop a Monte Carlo physics-informed neural networks (MC-PINNs) for solving the phonon Boltzmann transport equation (BTE) to model the multiscale heat conduction in solid materials. Specifically, we utilize a deep neural network to approximate the energy density function in BTE and encode the BTE as well as the boundary/initial conditions via the automatic differentiation, in MC-PINNs. In addition, we propose a novel two-step sampling approach to address the issues like “curse of dimensionality” and the inaccurate estimate of the equilibrium distribution functions in the existing sampling strategies that are widely used in PINNs. In particular, we first randomly sample a certain number of points in the temporal-spatial domains (Step I), and then we draw another number of random points in the solid angular domain (Step II). The residual points in MC-PINNs are constructed using the tensor product given the points generated at the above two steps. We conduct a series of numerical experiments to validate the accuracy of the MC-PINNs, including the steady quasi-1D/quasi-2D/3D as well as unsteady quasi-1D heat conduction problems spanning the diffusive and ballistic regimes. The results show the MC-PINNs is capable of achieving good accuracy for the multiscale heat conduction from 1D to 3D. Furthermore, the present method requires quite a small number of discrete points in the solid angular domain than in the deterministic methods for solving the BTE, especially for scenarios in the ballistic regimes. For instance, we only employ 64 points in the solid angular domain at each training step in MC-PINNs for Kn ranging from 0.1 to 10 in the 3D case. In contrast, we have $80\times 80$ discrete points in the solid angular domain of the numerical method to obtain converged results. As a result, the memory cost in the MC-PINN is about 1/6 of that in the numerical method. Further, we are able to achieve a similar computational efficiency for MC-PINNs to the numerical method in the 3D test case as we utilize the transfer learning technique in the MC-PINNs.
@@
We would like to discuss that the MC-PINN developed in this work is mesh-free in the temporal-spatial-angular domain, which is promising in real-world applications with complicated domains, for instance, the design of microdevices. In addition, the MC-PINN is capable of providing good accuracy for heat conduction ranging from diffusive to ballistic with only a small number of random discrete points in the solid angular domain at each training step. It can save lots of effort in numerical methods since the optimal discretization of the solid angular space at different regimes is generally different and should be carefully designed to ensure convergence in these methods. Also, the memory efficacy and the flexibility of treatment for the solid angular space in the present method make it a promising tool for modeling multiscale heat conduction in large-scale problems. Finally, we note that the MC-PINNs with the two-step sampling approach can be seamlessly extended to solve kinetic equations used in other disciplines, e.g., the Boltzmann-BGK model for multiscale flows, the radiation transfer equation for multiscale heat transfer, etc. We will leave these interesting topics as future studies.
@@ Appendix B Ablation study
To study the effect of the number of the points in the solid angular space at each training step on the predicted accuracy of MC-PINNs, we further conduct the following experiments based on two specific cases in Sec. 3.1, i.e., Kn $=1$ and $10$. In each case, we test 6 different numbers of the points in the solid angular space, i.e. $B_{\bm{s}}=2,4,8,16,32$ and 64 points at each training step. Also, we independently run MC-PINNs 5 times with different initializations for the hyperparameters to test the robustness of the MC-PINNs. The number of $B_{\bm{x}}$ is set to $40$, which is consistent with the case in Sec. 3.1. The predicted temperature from MC-PINNs is depicted in Fig. 15. It is observed that: (1) the computational errors in general decrease with the increase of $B_{\bm{s}}$, and (2) the relative $L_{2}$ errors between the results from MC-PINNs and the reference solutions in each case are below $1\%$ as $B_{\bm{s}}>2$. In other words, we are able to achieve good accuracy as we randomly sample more than 2 points in the solid angular domain at each training step in MC-PINNs for these two specific cases.
Figure 15: Quasi-1D film heat conduction: Relative $L_{2}$ errors of the temperature with different numbers of points in the solid angular space at each training step in PINNs. (a) $\rm{Kn}=1$, and (b) $\rm{Kn=10}$. Error bars in each figure represent the standard deviation calculated from 5 independent runs with different random seeds. The blue dashed line denotes the relative $L_{2}$ error for $1\%$.
We further present the loss histories for the representative case $\rm{Kn}=1.0$ with $B_{\bm{s}}=2$ and 16 in Fig. 16. As shown, the loss for MC-PINNs with 16 random points in the solid angular space after 50,000 training steps is about two orders smaller than that with 2 random points in the solid angular domain. Also, the fluctuations in the loss are smaller for the case with a larger number of points in the solid angular space at each training step.